# Inference-based GAN Video Generation

### v1

@@ Abstract
Video generation has seen remarkable progresses thanks to advancements in generative deep learning. Generated videos should not only display coherent and continuous movement but also meaningful movement in successions of scenes. Generating models such as Generative Adversarial Networks (GANs) or Variational Autoencoders (VAEs) and more recently Diffusion Networks have been used for generating short video sequences, usually of up to 16 frames. In this paper, we first propose a new type of video generator by enabling adversarial-based unconditional video generators with a variational encoder, akin to a VAE-GAN hybrid structure, in order to enable the generation process with inference capabilities. The proposed model, as in other video deep learning-based processing frameworks, incorporates two processing branches, one for content and another for movement. However, existing models struggle with the temporal scaling of the generated videos. In classical approaches when aiming to increase the generated video length, the resulting video quality degrades, particularly when considering generating significantly long sequences. To overcome this limitation, our research study extends the initially proposed VAE-GAN video generation model by employing a novel, memory-efficient approach to generate long videos composed of hundreds or thousands of frames ensuring their temporal continuity, consistency and dynamics. Our approach leverages a Markov chain framework with a recall mechanism, with each state representing a VAE-GAN short-length video generator. This setup allows for the sequential connection of generated video sub-sequences, enabling temporal dependencies, resulting in meaningful long video sequences.
@@ 2.1 Short-term Video Generation
2.2 Long-term Video Generation
@@
Recent studies in long-term video generation have managed to extend the number of generated frames by increasing the generated video length from 64 frames [tian2021mocoganhd] to 128 or 256 frames [wang2023styleinv, digan2022iclr], and even up to 1024 frames [ge2022tatsLong] or longer [nips2022dynamicLong, stylegan_v2022cvpr], while in earlier generative models the number of generated frames for videos have been of around 10 to 32 frames [Munoz_2021_tsgan, saito2017tgan, sun_twostreamvan_2020, tulyakov_mocogan_2018, vondrick2016vgan, wang_g3an_2020].
@@ 4 Long-term video generation
4.1 General long-temporal video generation framework
@@
Equations (10) and (11) are used to generate the initial and final parts of a video clip, where $\ominus$’ and $\oplus$’ represent the subtraction and addition, respectively, of frame differences representing movement. These frames are generated recursively for $j=1,\ldots,N$ video clips. By replacing Eq. (5) from EncGAN3 with Eq. (10) and (11) for the clip generation, we build the Markov chain, connecting each video clip to the next one. However, for building a long sequence, we should ensure that each generated video clip is smoothly connected to both its previous and next video clips by enforcing the continuity across the entire sequence, benefiting the modeling of temporal consistency in the long-term generated video. Note that when training with videos longer than 16 frames (e.g., extending to 20 frames), there is a dramatic increase in GPU memory requirements (as shown in Tab. 6), which directly limits the training video length. To overcome this, our proposed recall mechanism decomposes a long video into multiple temporally coherent sub-sequences, each requiring only a fixed amount of GPU memory. By connecting these sub-sequences using a Markov chain dependency on a designated reference frame, the overall memory usage remains constant regardless of the total video length, thereby enabling the generation of hundreds or even thousands of frames. Although this approach does not reduce the algorithmic time complexity or modifies the storage, it effectively circumvents in practice the GPU memory bottleneck. By enabling the temporal consistency in consecutive frames, we drop the frames generated in the second part by Eq. (11) for a video clip $j$, and replace them with the frames generated by Eq. (10) from the next video clip $j+1$. The exception is when adding the final video clip when these frames are considered as the ending part of the longer generated video $\widehat{\bf y}$. Eventually the long video is generated by concatenating $r$ frames from each video clip, generated according to Eq. (10), with the last $T_{c}-r$ frames being generated according to Eq. (11), as :
@@ 4.2 The recall mechanism
where we consider $N_{L}$ long-term videos, each split into $N_{C}$ overlapping clips, each containing $T_{c}$ frames, ${\bf x}_{m,i,j}$ represents an image frame while $\widehat{\bf x}_{m,i,j}$ is its reconstruction, $\widehat{\bf v}_{m,i,j}$ the reconstruction of the movement, as the difference between consecutive frames, associated with the frame $i$ form clip $j$ from the long-term video $m$. Meanwhile, $\{{\bf z}_{\bf x},{\bf z}_{\bf v}\}$ represent the latent spaces of the content and movement, modeled by the encoders $E_{c}$ and $E_{m}$, respectively. The reconstructions $\{\widehat{\bf x}_{ij}\}_{i=1}^{T_{c}}$ are made as close as possible to their corresponding original frames $\{{\bf x}_{ij}\}_{j=1}^{T_{c}}$ by the content and video reconstruction errors corresponding to the first and second terms in Eq. (15). The two KL divergence terms $D_{KL}$ ensure that the probabilities of the latent variable associated with the content ${\bf z}_{\bf x}$ and motion ${\bf z}_{\bf v}$, generated by the two encoders of parameters ${\theta}_{\bf x}$ and ${\theta}_{\bf v}$, are normal distributions, namely $p({\bf z}_{\bf x})$ and $p({\bf z}_{\bf v})$. By minimizing $L_{REncG}$, the two encoders implementing the variational distributions $q_{{\theta}_{\bf x}}({\bf z}_{\bf x}\,|\,{\bf x})$ and $q_{{\theta}_{\bf v}}({\bf z}_{\bf v}\,|\,{\bf v})$, encode the feature representations of content ${\bf z}_{\bf x}$ and motion ${\bf z}_{\bf v}$, for enabling a better data reconstruction by the Generator.
@@ 5.2.3 Ablation Studies
EncGAN3 module components. We test the contribution of the components to the efficiency of the EncGAN3, such as the Encoder and the three-stream processing pipeline, F-SA module in the Generator $G$, considering three datasets for the training, UvA, Weizmann and KTH, with the results provided in Table 4. We consider assessing the results when removing the Encoder (Enc) or the F-SA processing modules, and also when removing the auxiliary reconstruction streams for the content $G_{C}$ and movement $G_{T}$. From the results provided in Table 4, we observe that the presence of two auxiliary streams $G_{C}$ and $G_{T}$, as well as the Encoder and F-SA module, are important for the performance of EncGAN3.
@@
Video-sampling strategies for training EncGAN3. We provide the results when considering two different sampling strategies when training EncGAN3 on the UvA database : step sampling and uniform sampling. The step sampling used in G3AN [wang_g3an_2020], consists of randomly selecting a starting frame and then sampling the following video frames with a sampling step of 2 or 3 for the entire number of frames considered. Meanwhile, we consider the uniform sampling that divides a video clip into 16 sets with equal numbers of frames, and then randomly sample one of the frames from within each set. The uniform sampling represent the temporal information covering an entire video sequence. The uniform sampling models better the variation of the information over the entire video sequence when compared to the step sampling. When considering uniform sampling, the frames from videos generated by EncGAN3 trained on the uniformly sampled training set, after 100 epochs, and after 5000 epochs, are provided on the top and bottom row of Figure 7(a), respectively. Meanwhile, frames from videos generated by G3AN, after being trained on the uniformly sampled training set for 100 and 5000 epochs are shown at the top and bottom rows of Figure 7(b), respectively. From these results based on the more complex uniformly sampled training set, it can be observed that EncGAN3 provides significantly better results than G3AN given that by using an Encoder for inference leads to more stable modeling, quicker convergence and better generation results.
@@ 5.3 Long-temporal generated video sequences results
In this section, we present and analyze the long-temporal video generation results, when using REncGAN3, described in Section 4.2, which employs the recall mechanism, for generating video sequences of hundreds of frames. We consider training on the Tai-Chi-HD [Taichi] and Sky-Timelapse [Sky] databases for generating videos with over a hundred frames by following the long-term video generation methodology. Taichi-HD (Taichi), contains 264 in-the-wild Tai-Chi videos from YouTube showing different practitioners performing complex movement sequences with different backgrounds, while the Sky-timelapse (Sky) dataset contains videos displaying movements of clouds in the sky, under various lighting conditions. Taichi dataset contains just 4 videos of lengths over 1024 frames while Sky contains more than 200 such long-term videos. REncGAN3 is implemented using the same settings as EncGAN3, which are described in the first paragraph from Section 5. The generation of $128\times 128$ pixels videos trained on Taichi and Sky datasets are created using one A60 GPU with 46GB memory using the Ubuntu operating system.
@@ 5.3.1 Qualitative results for the long-temporal generated videos
In Figure 11, the frames generated by REncGAN3 show clouds gradually covering the trees in the left panel and with less noise than those in the second row which also display some unrealistic artifacts. Moreover, REncGAN3 is able to generate videos of various lengths without using interpolation or extrapolation, as DIGAN does in [digan2022iclr], while it relies on a simple recall mechanism, which connects short-term clips, as proposed in this paper. In addition, as shown in Figure 14, although generating long videos of the human action Taichi sequences with good motion dynamics is hard [digan2022iclr] due to the strict constraints of physical body specific to these actions, our REncGAN3 model provides visual results of good quality when generating the video sequence from sampled latent codes, as shown in Figure 14 (b), as well as by using the mean value, as shown in Figure 14 (a). The obvious motion dynamics benefits from the mixture of the clip-by-clip and frame-by-frame modeling together with the generation processes in the recall mechanism, as proposed in this paper.
@@
5.3.2 Quantitative results for the long-temporal video generation
@@
To further measure the quality degradation, we evaluate the consistency and continuity over the generated long video sequences by segmenting them into consecutive non-overlapping video segments of 16 frames from the long-term generated video, and then we evaluate the FVD and the Fréchet Inception Distance (FID) [FID] in the video domain [wang_g3an_2020], on each video segment individually. Similar to FVD, lower video FID values means better visual quality and spatial-temporal consistency of generated videos. In this way, we track the evaluation of the entire video’s continuity and consistency in the content and movement quality. Such an evaluation approach can be applied to videos of arbitrary lengths. In Figure 12, we provide the video FID values for three long-term generated videos after training on the Taichi dataset as well as their average. Two of the generated videos display consistency in good FID scores, while the one indicated in red and labeled as 560f, displaying more complex variation with some segments characterized by high FID scores. The recall mechanism in REncGAN3 merges short-term video clips instead of generating them frame by frame, displaying good quality consistency and addressing the degradation problem present in the long-term videos generated by other methods, as shown by the FID results from Figure 12.
@@ 5.4 Comparative evaluation of long video generation models
Since most videos in the datasets used for training are of approximately 100 frames long, and many methods that simply increase the latent code size struggle to maintain good frame quality beyond this range [digan2022iclr, wang2023styleinv], the primary comparison between EncGAN3 and REncGAN3 focuses on generating videos of around 100 frames. However, we consider generating videos of more than 400 frames for REncGAN3 and LEncGAN3, given that both are designed specifically for long video generation. The results for REncGAN3 and LEncGAN3 are provided in Figure 16(a) and 16(b) when considering training on the TaiChi database. From the frames shown in these figures, it can be observed that LEncGAN3 tends to produce artifacts in the background, whereas REncGAN3 consistently delivers higher-quality videos, particularly excelling in the generation of facial features and by smoothly modeling the movement of the hands. These results highlight the superior performance of REncGAN3 in long video generation tasks.

### v2

@@ Abstract
Video generation has seen remarkable progress thanks to advancements in generative deep learning. However, generating long sequences remains a significant challenge. Generated videos should not only display coherent and continuous movement but also meaningful movement in successions of scenes. Models such as GANs, VAEs, and Diffusion Networks have been used for generating short video sequences, typically up to 16 frames. In this paper, we first propose a new type of video generator by enabling adversarial-based unconditional video generators with a variational encoder, akin to a VAE-GAN hybrid structure. The proposed model, as in other video deep learning-based processing frameworks, incorporates two processing branches, one for content and another for movement. However, existing models struggle with the temporal scaling of the generated videos. Classical approaches often result in degraded video quality when attempting to increase the generated video length, especially for significantly long sequences. To overcome this limitation, our research study extends the initially proposed VAE-GAN video generation model by employing a novel, memory-efficient approach to generate long videos composed of hundreds or thousands of frames ensuring their temporal continuity, consistency and dynamics. Our approach leverages a Markov chain framework with a recall mechanism, where each state represents a short-length VAE-GAN video generator. This setup enables the sequential connection of generated video sub-sequences, maintaining temporal dependencies and resulting in meaningful long video sequences.
Keywords: Video generation, Generative Adversarial Networks, Variational Autoencoder, VAE-GAN, long-video generation, Markov chain.
@@ 2.1 Short-term Video Generation
2.2 Existing long-term Video generation methods
@@
Recent studies in long-term video generation have managed to extend the number of generated frames by increasing the generated video length from 64 frames [tian2021mocoganhd] to 128 or 256 frames [wang2023styleinv, digan2022iclr], and even up to 1024 frames [ge2022tatsLong] or longer [nips2022dynamicLong, stylegan_v2022cvpr], while in earlier generative models the number of generated frames for videos has been of around 10 to 32 frames [Munoz_2021_tsgan, saito2017tgan, sun_twostreamvan_2020, tulyakov_mocogan_2018, vondrick2016vgan, wang_g3an_2020].
@@ 4 Long-term Video Generation
4.1 General framework for long-temporal video generation
@@
Equations (10) and (11) are used to generate the initial and final parts of a video clip, where $\ominus$’ and $\oplus$’ represent the subtraction and addition, respectively, of frame differences representing movement. These frames are generated recursively for $j=1,\ldots,N$ video clips. By replacing Eq. (5) from EncGAN3 with Eq. (10) and (11) for the clip generation, we build the Markov chain, connecting each video clip to the next one. However, for building a long sequence, we should ensure that each generated video clip is smoothly connected to both its previous and next video clips by enforcing the continuity across the entire sequence, benefiting the modeling of temporal consistency in the long-term generated video. Note that when training with videos longer than 16 frames (e.g., extending to 20 frames), there is a dramatic increase in GPU memory requirements (as shown in Tab. 6), which directly limits the training video length. To overcome this, our proposed recall mechanism decomposes a long video into multiple temporally coherent sub-sequences, each requiring only a fixed amount of GPU memory. By connecting these sub-sequences using a Markov chain dependency on a designated reference frame, the overall memory usage remains constant regardless of the total video length, thereby enabling the generation of hundreds or even thousands of frames. Although this approach does not reduce the algorithmic time complexity or modify the storage, it effectively circumvents in practice the GPU memory bottleneck. By enabling the temporal consistency in consecutive frames, we drop the frames generated in the second part by Eq. (11) for a video clip $j$, and replace them with the frames generated by Eq. (10) from the next video clip $j+1$. The exception is when adding the final video clip when these frames are considered as the ending part of the longer generated video $\widehat{\bf y}$. Eventually the long video is generated by concatenating $r$ frames from each video clip, generated according to Eq. (10), with the last $T_{c}-r$ frames being generated according to Eq. (11), as :
@@ 4.2 The recall mechanism
where we consider $N_{L}$ long-term videos, each split into $N_{C}$ overlapping clips, each containing $T_{c}$ frames, ${\bf x}_{m,i,j}$ represents an image frame while $\widehat{\bf x}_{m,i,j}$ is its reconstruction, $\widehat{\bf v}_{m,i,j}$ the reconstruction of the movement, as the difference between consecutive frames, associated with the frame $i$ from clip $j$ from the long-term video $m$. Meanwhile, $\{{\bf z}_{\bf x},{\bf z}_{\bf v}\}$ represent the latent spaces of the content and movement, modeled by the encoders $E_{c}$ and $E_{m}$, respectively. The reconstructions $\{\widehat{\bf x}_{ij}\}_{i=1}^{T_{c}}$ are made as close as possible to their corresponding original frames $\{{\bf x}_{ij}\}_{j=1}^{T_{c}}$ by the content and video reconstruction errors corresponding to the first and second terms in Eq. (15). The two KL divergence terms $D_{KL}$ ensure that the probabilities of the latent variable associated with the content ${\bf z}_{\bf x}$ and motion ${\bf z}_{\bf v}$, generated by the two encoders of parameters ${\theta}_{\bf x}$ and ${\theta}_{\bf v}$, are normal distributions, namely $p({\bf z}_{\bf x})$ and $p({\bf z}_{\bf v})$. By minimizing $L_{REncG}$, the two encoders implementing the variational distributions $q_{{\theta}_{\bf x}}({\bf z}_{\bf x}\,|\,{\bf x})$ and $q_{{\theta}_{\bf v}}({\bf z}_{\bf v}\,|\,{\bf v})$, encode the feature representations of content ${\bf z}_{\bf x}$ and motion ${\bf z}_{\bf v}$, for enabling a better data reconstruction by the Generator.
@@ 5.2.3 Ablation Studies
EncGAN3 module components. We test the contribution of the components to the efficiency of the EncGAN3, such as the Encoder and the three-stream processing pipeline, F-SA module in the Generator $G$, considering three datasets for the training, UvA, Weizmann and KTH, with the results provided in Table 4. We consider assessing the results when removing the Encoder (Enc) or the F-SA processing modules, and also when removing the auxiliary reconstruction streams for the content $G_{C}$ and movement $G_{T}$. From the results provided in Table 4, we observe that the presence of two auxiliary streams $G_{C}$ and $G_{T}$, as well as the Encoder and F-SA module, is important for the performance of EncGAN3.
@@
Video-sampling strategies for training EncGAN3. We provide the results when considering two different sampling strategies when training EncGAN3 on the UvA database : step sampling and uniform sampling. The step sampling used in G3AN [wang_g3an_2020], consists of randomly selecting a starting frame and then sampling the following video frames with a sampling step of 2 or 3 for the entire number of frames considered. Meanwhile, we consider the uniform sampling that divides a video clip into 16 sets with equal numbers of frames, and then randomly sample one of the frames from within each set. The uniform sampling represents the temporal information covering an entire video sequence. The uniform sampling models better the variation of the information over the entire video sequence when compared to the step sampling. When considering uniform sampling, the frames from videos generated by EncGAN3 trained on the uniformly sampled training set, after 100 epochs, and after 5000 epochs, are provided on the top and bottom row of Figure 7(a), respectively. Meanwhile, frames from videos generated by G3AN, after being trained on the uniformly sampled training set for 100 and 5000 epochs are shown at the top and bottom rows of Figure 7(b), respectively. From these results based on the more complex uniformly sampled training set, it can be observed that EncGAN3 provides significantly better results than G3AN given that by using an Encoder for inference leads to more stable modeling, quicker convergence and better generation results.
@@ 5.3 Long-temporal generated video sequences results
In this section, we present and analyze the long-temporal video generation results, when using REncGAN3, described in Section 4.2, which employs the recall mechanism, for generating video sequences of hundreds of frames. We consider training on the Tai-Chi-HD [Taichi] and Sky-Timelapse [Sky] databases for generating videos with over a hundred frames by following the long-term video generation methodology. Taichi-HD (Taichi), contains 264 in-the-wild Tai-Chi videos from YouTube showing different practitioners performing complex movement sequences with different backgrounds, while the Sky-timelapse (Sky) dataset contains videos displaying movements of clouds in the sky, under various lighting conditions. Taichi dataset contains just 4 videos of lengths over 1024 frames while Sky contains more than 200 such long-term videos. REncGAN3 is implemented using the same settings as EncGAN3, which are described in the first paragraph from Section 5. The generation of $128\times 128$ pixels videos trained on Taichi and Sky datasets is created using one A60 GPU with 46GB memory using the Ubuntu operating system.
@@ 5.3.1 Qualitative results for the long-temporal generated videos
In Figure 11, the frames generated by REncGAN3 show clouds gradually covering the trees in the left panel and with less noise than those in the second row which also display some unrealistic artifacts. Moreover, REncGAN3 is able to generate videos of various lengths without using interpolation or extrapolation, as DIGAN does in [digan2022iclr], while it relies on a simple recall mechanism, which connects short-term clips, as proposed in this paper. In addition, as shown in Figure 14, although generating long videos of the human action Taichi sequences with good motion dynamics is hard [digan2022iclr] due to the strict constraints of physical body specific to these actions, our REncGAN3 model provides visual results of good quality when generating the video sequence from sampled latent codes, as shown in Figure 14 (b), as well as by using the mean value, as shown in Figure 14 (a). The obvious motion dynamics benefit from the mixture of the clip-by-clip and frame-by-frame modeling together with the generation processes in the recall mechanism, as proposed in this paper.
@@
5.3.2 Quantitative results for long-temporal video generation
@@
To further measure the quality degradation, we evaluate the consistency and continuity over the generated long video sequences by segmenting them into consecutive non-overlapping video segments of 16 frames from the long-term generated video, and then we evaluate the FVD and the Fréchet Inception Distance (FID) [FID] in the video domain [wang_g3an_2020], on each video segment individually. Similar to FVD, lower video FID values mean better visual quality and spatial-temporal consistency of generated videos. In this way, we track the evaluation of the entire video’s continuity and consistency in the content and movement quality. Such an evaluation approach can be applied to videos of arbitrary lengths. In Figure 12, we provide the video FID values for three long-term generated videos after training on the TaiChi dataset as well as their average. Two of the generated videos display consistency in good FID scores, while the one indicated in red and labeled as 560f, displays more complex variation with some segments characterized by high FID scores. The recall mechanism in REncGAN3 merges short-term video clips instead of generating them frame by frame, displaying good quality consistency and addressing the degradation problem present in the long-term videos generated by other methods, as shown by the FID results from Figure 12.
@@ 5.4 Comparative evaluation of long video generation models
Since most videos in the datasets used for training are of approximately 100 frames long, and many methods that simply increase the latent code size struggle to maintain good frame quality beyond this range [digan2022iclr, wang2023styleinv], the primary comparison between EncGAN3 and REncGAN3 focuses on generating videos of around 100 frames. However, we consider generating videos of more than 400 frames for REncGAN3 and LEncGAN3, given that both are designed specifically for long video generation. The results for REncGAN3 and LEncGAN3 are provided in Figures 16(a) and 16(b) when considering training on the TaiChi database. From the frames shown in these figures, it can be observed that LEncGAN3 tends to produce artifacts in the background, whereas REncGAN3 consistently delivers higher-quality videos, particularly excelling in the generation of facial features and by smoothly modeling the movement of the hands. These results highlight the superior performance of REncGAN3 in long video generation tasks.